### v1

# Federated Continual Learning: Concepts, Challenges, and Solutions

Abstract
Federated Continual Learning (FCL) has emerged as a robust solution for collaborative model training in dynamic environments, where data samples are continuously generated and distributed across multiple devices. This survey provides a comprehensive review of FCL, focusing on key challenges such as heterogeneity, model stability, communication overhead, and privacy preservation. We explore various forms of heterogeneity and their impact on model performance. Solutions to non-IID data, resource-constrained platforms, and personalized learning are reviewed in an effort to show the complexities of handling heterogeneous data distributions. Next, we review techniques for ensuring model stability and avoiding catastrophic forgetting, which are critical in non-stationary environments. Privacy-preserving techniques are another aspect of FCL that have been reviewed in this work. This survey has integrated insights from federated learning and continual learning to present strategies for improving the efficacy and scalability of FCL systems, making it applicable to a wide range of real-world scenarios.
keywords:
Federated continual learning , Continual learning , Federated learning , Incremental learning , Non-stationary data , Concept drift
journal: Computer Science Review
[…]
[label1]organization=Faculty of Computer Science, University of New Brunswick, addressline=550 Windsor Street, city=Fredericton, postcode=E3B 5A3, state=NB, country=Canada
[…]
[label2]organization=Department of Electrical and Computer Engineering, University of Windsor, addressline=401 Sunset Ave, city=Windsor, postcode=N9B 3P4, state=ON, country=Canada
1 Introduction
The increasing number of interconnected devices and the growing demand for the use of intelligent systems in practice has sparked the interest into the concepts of distributed and incremental learning [1, 2]. A noticeable trend is Federated Learning (FL), which has established itself as an efficient mechanism to accomplish collaborative training of machine learning models across decentralized devices and resolve privacy and communication issues by keeping data where it belongs [3, 4]. On the other hand, Continual Learning (CL) aims to solve the problem of forcing models to learn and forget knowledge as the target data distribution changes, and, thus, it is relevant to non-static scenarios [1, 5]. There is growing interest in the broader distribution and generation of information. Consequently, FCL has attracted significant attention from researchers for its ability to integrate the strengths of FL and CL, making it ideal for practical applications such as distributed and incremental learning.
FCL draws the foundation from the literature in FL and CL [6]. FL addresses key challenges such as data and system heterogeneity, communication overhead, and privacy preservation in distributed systems. It enables global model training across diverse clients without centralizing data. However, it assumes static data distributions, making it less suited for dynamic environments where data streams continuously. While CL addresses issues with the presence of non-stationary data, a method that could deal with issues through techniques such as experience replay, regularization, or modification in architecture so as to ensure models stay time-robust and adapt to new incoming tasks without forgetting obtained knowledge in the previous state. FCL bridges between these domains by integrating FL concepts along with CL to effectively control evolving data [7, 8]. For example, heterogeneity and privacy challenges brought about by FL are more complex in FCL because data streams dynamically, while solutions provided for incremental learning should be adapted to work under FL constraints such as non-IID data and limited communication [9]. These cross-domain interactions help us to better understand how existing research on FL and CL is fundamentally important in advancing FCL.
FCL is imperative in facilitating intelligent systems, where data is both distributed and non-stationary. The application of FCL ranges from personalized healthcare to real-time IoT analytics and edge-based autonomous systems [10, 11]. Some of the unique challenges in FCL, such as addressing global and local catastrophic forgetting, heterogeneous client behavior, and scalable knowledge transfer across clients, remain barely explored in the literature. These challenges require new methodologies that effectively combine FL’s focus on collaboration and privacy with CL’s ability to adapt over time. Furthermore, practical issues such as dynamic client participation, resource constraints, and real-world non-IID data distributions make things even more complicated. Despite its promise, the area of FCL lacks a comprehensive overview of its challenges, methodologies, and applications. This gap calls for a dedicated survey to consolidate knowledge, identify open problems, solutions, and guide future research.
Currently, the literature offers only a couple of surveys on FCL. The first published survey categorizes methods based on knowledge fusion mechanisms and introduces synchronous and asynchronous FCL frameworks[6]. In addition, a recent preprint focuses on FCL within Edge-AI environments, with an emphasis on task-specific classifications and hardware-algorithm co-design [12].While these surveys have contributed valuable insights on FCL, focusing on knowledge fusion, task-specific classifications, and Edge-AI contexts. However, they lack a comprehensive taxonomy of key challenges, such as heterogeneity, communication overhead, and model stability, and offer limited cross-domain insights from FL and CL. This survey aims to build on prior works to address these gaps.
This survey addresses the pressing need for a detailed examination of FCL, motivated by its relevance to real-world applications and the open problems in the field. By systematically exploring how FL and CL challenges converge in FCL, this work aims to provide researchers with a clear understanding of the domain’s current state and future directions. Specifically, it highlights critical issues such as catastrophic forgetting in federated settings, non-stationary client environments, and the interplay of privacy, communication, and computational constraints. By bridging the gap between FL and CL, this survey aims to serve as a foundational reference for developing robust and scalable FCL systems, ultimately contributing to the broader adoption of intelligent, adaptive, and decentralized AI solutions.
The remainder of this survey is organized as follows. Section 2 reviews the preliminaries of concepts such as FL, CL, and FCL. Section 3 discusses challenges and solutions associated with FL comonents of FCL. Section 4 elaborates on difficaulties of CL and the potential solutions. Section 5 explains problems that are specific to FCL and overviews the key approaches for mitigating them. Section 6 elaborates on future research directions that are crucial in FCL. Finally, the paper is concluded in Section 7.
2 Federated Continual Learning
The data-driven world requires learning from continuous data in a distributed manner, while respecting privacy. FCL emerges as a robust solution for scenarios involving both collaborative learning and streaming data. FCL seamlessly integrates FL principles with the capability to manage continuous data streams, offering a versatile and efficient approach to contemporary data challenges. By leveraging FCL, it is possible to achieve the dual objectives of collaboratively training models without centralized data collection and maintaining up-to-date models through continual learning from streaming data.
The first step toward understanding FCL is to understand its base ideas: FL and CL. FL refers to a variety of model training that involves multiple devices/nodes in collaboration, not sharing raw data with each other for privacy and scalability reasons. On the other hand, CL deals with models’ capabilities of incremental learning on non-IID evolving data streams, while retaining knowledge of previous tasks. These two paradigms, while different, meet at the intersection of FCL in the quest to build models that continuously learn from new data without violating the privacy and heterogeneity of decentralized systems. In-depth knowledge of FL and CL puts one in a better perspective to understand the challenges and opportunities that come with FCL.
2.1 Federated Learning
FL is a collaborative method for training machine learning models across multiple devices or nodes with the guarantee that raw data would remain on the local devices. It is motivated by the urgent need for privacy-preserving and efficient learning systems in recent distributed environments, such as smartphones, IoT devices, and edge computing networks. In particular, unlike traditional, top-down approaches to learning in a decentralized system, FL lets owners of data participate in training the model without necessarily revealing personally identifiable or commercially sensitive information.
By keeping data decentralized, FL protects user privacy and also allows compliance with strict data protection regulations, such as GDPR. It reduces the risk of data breaches and ensures that the training data remains local; hence, trust among participants can be fostered. FL also leverages the heavy computational power of distributed devices, hence enhancing scalability, and, thus, enabling the creation of models that can make the most out of diverse data sources. This generally results in more generalizable and robust models for applications in which data across users may vary substantially, such as personalized recommendation systems, healthcare, and autonomous systems. FL will be capable of harnessing decentralized data and resources, thus becoming one of the cornerstones for developing intelligent, privacy-aware systems in modern, connected environments.
2.2 Continual Learning
In dynamic environments such as real-time sensor networks or streaming applications, data is continuously generated and needs to be processed in a timely manner. For instance, consider a smart city scenario where traffic cameras, weather sensors, and social media feeds constantly produce data streams. In such an environment,which is known as non stationary environment, it is essential for models to be updated continuously with new data to maintain their performance. However, traditional models often struggle to adapt to changes in data distributions, leading to a decrease in performance over time. This challenge is addressed by CL, which focuses on the ability to incrementally learn from new data, while retaining knowledge from previous tasks [1, 13]. By effectively managing the evolving nature of data, CL ensures that models remain robust and adaptable, preventing the degradation of performance that can occur with static training methods.
CL provides several significant advantages that enhance the effectiveness and adaptability of machine learning models. These key benefits include incremental learning, knowledge transfer, avoiding catastrophic forgetting, adaptability and flexibility, resource efficiency, improved generalization, and application versatility. Firstly, CL allows models to learn incrementally, building on previously acquired knowledge and enabling adaptation to new tasks without forgetting old ones, which is particularly useful in dynamic environments, where data evolves over time [5]. Secondly, it facilitates the transfer of knowledge from one task to another, improving learning efficiency for new tasks by leveraging previously acquired skills, thereby leading to faster learning and better performance on related tasks [14, 15].
Furthermore, a main advantage is its ability to mitigate catastrophic forgetting, using techniques such as experience replay, regularization, and dynamic architecture growth to maintain old knowledge [16]. Additionally, CL models are highly adaptable, making them suitable for non-stationary, continuously changing real-world applications [17]. Moreover, they emphasize resource efficiency by reusing existing knowledge and minimizing the need for retraining, which leads to significant savings in computational resources and time. For instance, [18] introduces a method for performing text classification using pre-trained transformers on a sequence of classification tasks, achieving significant memory efficiency and faster inference time compared to state-of-the-art methods, while maintaining competitive predictive performance without requiring extensive retraining. By continually learning from new data, CL models also improve their generalization capabilities, making them more robust to data variations and better at handling unseen scenarios [19]. Finally, CL has broad applications across various domains, including robotics, healthcare, and autonomous systems, where the ability to learn and adapt continuously is critical [20].
3 Challenges of Federated Learning
Part of the challenges in FCL are assocciated to the FL structure of this approach as detailed in this section.
3.1 Heterogeneity
In collaborative learning, different clients contribute to training, which can enhance overall training performance by leveraging server computational power and improving convergence rates and globalization [21, 22]. However, this diversity can introduce heterogeneity, stemming from factors such as geographical differences, variations in data source distribution, differing data formats, inconsistent data quality, variations in data collection methods, discrepancies in labeling practices, and differences in feature representation [23]. Heterogeneity, which refers to the presence of differences and diversity within a set of elements, is one of the prominent challenges in collaborative learning. [24] highlights the severe impacts of device and behavioral heterogeneity on model performance, showing substantial degradation in model quality and fairness. Another study on dialogue generation focuses on branch homogeneity problems in collaborative learning and proposes methods to enhance branch heterogeneity for improved performance [25]. Tailored strategies are essential for managing heterogeneity, as it can hinder the effectiveness of collaborative learning. Effective management is crucial to fully realize the potential of collaborative learning. This subsection explains different types of heterogeneity and proposes methods to address them effectively, categorizing heterogeneity into two primary types: data heterogeneity and system heterogeneity.
Data heterogeneity arises from the varied nature of data contributed by different clients. Figure 1 delineates the subsets of data heterogeneity. Table 1 highlights datasets demonstrating such heterogeneity, including non-IID distributions and imbalanced data. Understanding data heterogeneity is crucial for effective model training and integration [26, 27]. Data heterogeneity is generally classified into statistical and conceptual subsets, each of which presents a distinct challenge that needs to be addressed in detail.
{forest}
[…]
Heterogeneity [Statistical
Heterogeneity [Non-IID Data [Changes in
Input Space] [Changes in
Behaviour] ] [Imbalanced Data [Unequal
Sample
Sizes] [Skewed Data
Distribution] ] ] [Conceptual
Heterogeneity [Label Distribution
Heterogeneity [Diverse
Label Sets] [Inconsistent
Labeling] ] [Feature Distribution
Heterogeneity [Varying
Feature Sets] [Feature
Scaling
Differences] ] ] ]
Figure 1: Sub-categories of data heterogeneity in collaborative learning.
3.1.1 Statistical Heterogeneity
Statistical heterogeneity refers to differences in the statistical properties of datasets used by different nodes in a distributed learning system. This can include variations in data distributions, sample sizes, and data quality across different sources. Two critical subsets of statistical heterogeneity are Non-IID (Independent and Identically Distributed) data and imbalanced data.
Non-IID Data
Many machine learning algorithms assume that the training data are independent and identically distributed (IID), meaning each data point is independent of others and follows the same probability distribution. When data from different clients violate this assumption, they are considered non-IID. These dependencies arise from differences in input space or behavioral differences among clients.
Table 1: Examples of data heterogeneity considered in FCL research. Data Format: categorical (C), numerical (N), and mixed (M). Research field: computer vision (CV), healthcare (Health), cybersecurity (Cyber), Internet of Things (IoT), natural language processing (NLP), and entertainment (EN). Task type: instance incremental (Inst), domain incremental (Dom), task incremental (Task), feature incremental (Feat), and class incremental (Cls). Output type: label (Label), score (S), and partition (Partition). ✗ and * denote not applicable and not mentioned.
Approaches Data Format Distribution Drift Type Task Type Domain Type
Fed-ReMECS [28] C Non-IID * Inst CV Label
ISCN [29] N Non-IID * Inst/Feat Cyber Label
[30] C Non-IID * Cls/Task CV Label
FedSpace [8] C Non-IID R Task CV Label
FedConD [31] C Non-IID R Dom CV Label
ASO-Fed [32] C Non-IID * Inst/Dom CV Label
[33] M Non-IID V Inst EN/NLP Label
CHFL [7] N/M Non-IID * Dom/Feat CV/IoT Label
Cross-FCL [34] C Non-IID * Task CV Label
FedDrift-Eager/FedDrift [35] N Non-IID H Dom Label
Master-FL [36] N/M Non-IID R/V Dom CV Label
SOFed [37] N Non-IID V Inst/Dom CV Partition
FLARE [38] N Non-IID H Dom IoT/CV Label
[39] N Non-IID R Task/Cls CV Label
FairFedDrift [40] N * R Dom CV Label
[41] N IID/Non-IID R/V Inst CV Label
ADMM-FedMeta [42] N Non-IID H Task/Cls CV Label
FedINC [43] N Non-IID H Inst/Task/Cls CV Label
Adaptive-FedAVG [44] N IID H Task/Cls CV Label
AFAFed [11] N IID+Non-IID R Inst/Dom/Task IoT Label
O-GFML/PSO-GFML [45] N Non-IID * Task Score
GLFC [46] N Non-IID R Task/Cls CV Label
[47] N IID/Non-IID R Inst CV Label
CoOptFL/DYNAMITE [48] N IID/Non-IID R/V/H Dom/Task/Inst CV Label
Fed-IW&DS [49] N Non-IID R/V/H Dom CV Label
[50] M Non-IID H Dom Cyber Label
FedSKF [51] N IID/Non-IID R Cls CV Label
[52] N IID R Dom CV Label
[53] N Non-IID * Inst CV Label
FCIL-MSN [54] N Non-IID H Task/Cls CV Label
FedADC [55] N Non-IID H Dom CV Label
FCL4SR [56] M * H Dom CV/IoT Label
[57] N * * Task/Cls Health/IoT Label
FCL-SBLS [58] M * * Task Cyber Score
FedStream [59] C/N/M IID/Non-IID V * CV/Cyber/IoT Label
FedStream [60] N/M * R/V/H Inst/Dom Cyber/IoT Label
[61] N Non-IID V/H Dom/Task/Cls CV Label
FedProK [62] N IID/Non-IID R/H Task/Feat/Cls CV Label
FedKNOW [63] N Non-IID R/V/H Task CV Label
FedRCIL [64] N IID/Non-IID R/V/H Task CV Label
LCFIL [65] N IID/Non-IID R/V/H Task/Cls CV Label
Re-Fed [66] N IID/Non-IID R/V/H Dom/Task/Cls CV Label
ICMFed [67] N Non-IID V Task CV Label
LGA [68] N Non-IID H Task/Cls CV Label
[69] N IID/Non-IID V Inst CV/IoT Label
Flash [70] N Non-IID R/H Task CV/NLP Label
[71] M * * Task NLP Label
FedPC [72] N Non-IID V * CV Label
cTD-$\alpha$MAML[73] N IID/Non-IID R/V/H Task CV/Health Label
SFLEDS [74] N/M * R/V/H Inst/Task/Cls Cyber/IoT Label
FedNN [75] N Non-IID R/H Dom CV Label
Fed-SOINN [76] M IID/Non-IID * Task/Cls Cyber Label
RRA-FFSCIL [77] N Non-IID H Task/Cls CV Label
[78] N Non-IID R Inst CV Label
[79] N/M Non-IID * Task CV/IoT Label
[80] N IID H Dom/Task Label
FL-IIDS [81] M IID R Task/Cls Cyber Label
FedViT [82] N Non-IID V/H Task CV Label
CFeD [83] N IID/Non-IID R/V/H Dom/Task/Cls CV/NLP Label
FedET [84] N Non-IID V/H Task/Cls CV/NLP Label
FedWeIT [85] N Non-IID V Task CV Label
[86] N * * * NLP Label
[87] N Non-IID R/H Cls CV Label
[88] N * * Task IoT Label
SLMFed [89] N Non-IID * * CV Label
TARGET [90] N IID/Non-IID R/V/H Task CV Label
HFIN [91] M Non-IID R/V/H Task/Cls Cyber Score
[92] N Non-IID * * CV Label
Differences in input space occur when client $i$ and client $j$ contribute to the training and their local datasets have different distributions $P(x_{i})\neq P(x_{j})$. On the other hand, behavioral differences refer to variations in the way different clients generate, collect, or label their data, leading to different perspectives on events for the machine learning models $P(y_{i}|x_{i})\neq P(y_{j}|x_{j})$. These differences can impact the performance of machine learning models trained on such datasets. Imagine a smart home energy management system deployed across households in different regions, each acting as a client contributing data to a global energy optimization model. Client $i$ in Region A (temperate climate) has energy usage peaks during winter for heating and summer for cooling, with high Heating, Ventilation, and Air Conditioning (HVAC) consumption. Client $j$ in Region B (tropical climate) has consistently high energy usage for cooling year-round. Thus, their energy consumption patterns differ significantly ($P(x_{i})\neq P(x_{j})$). Households in Region A follow consistent daily routines, leading to predictable energy usage peaks in the morning and evening. In contrast, households in Region B have varied routines, with unpredictable energy usage patterns throughout the day. These behavioral differences lead to variations in energy consumption metrics $P(y_{i}|x_{i})\neq P(y_{j}|x_{j})$. To address Non-IID data as a prominent challenge in distributed systems, the proposed methods are reviewed here to clarify the importance of each way.
To address changes in the input space through clients, three different methods are proposed. Firstly, domain transformation involves techniques designed to align the data distributions from multiple clients to create a more unified input space for a machine learning model. One approach focuses on domains with particular features, leveraging the unique characteristics of each domain to enhance performance. For instance, studies [93, 94, 95] address the issue that each domain may have its own unique set of features to characterize samples, leading to incompatibilities across domains, and they develop methods to extract a common feature representation. To talk in detail, deep learning models are utilized for learning complex relations between features [96], and domain generalization [97]. Moreover, [98] proposed an approach in which feature transformation and KL divergence minimization are used to align source and target data distributions. Another approach, performed in [99, 100, 101, 102, 103, 104], is domain factorization, which decomposes data into shared and domain-specific components using methods such as matrix factorization. This isolates common patterns and unique features, allowing models to better understand the underlying structure. These techniques improve the robustness and effectiveness of machine learning models across diverse datasets.
Moreover, due to the significant variations in data distributions across different clients, personalization in FL is essential for addressing changes in the input space. This approach involves tailoring a global machine learning model to better fit the unique data and usage patterns of individual users. By doing so, the model becomes not only broadly effective but also finely tuned to the specific needs and behaviors of each user. [105] underscores the importance of personalization, while [106] explores methods to enhance the performance of FL models by incorporating personalization techniques, thereby addressing the challenges posed by data heterogeneity across different clients. Personalization methods are categorized into two subsets of client level personalization and group level personalization. Client level personalization tailors the model to the specific data characteristics and patterns of each individual client. For instance, [107] address the heterogeneity of data across different clients by learning personalized self-attention layers for each client, while aggregating other parameters among clients. Also, [108] aims to customize the global machine learning model to better fit the unique data distributions of individual clients. Another idea, explored in [109, 110, 111], involves adapting algorithms from the Model-Agnostic Meta-Learning (MAML) framework for use in FL settings. Instead of personalizing the model for each individual client, in group level personalization the model is tailored for groups with homogeneous data distributions. [112] leverages group-level personalization by clustering clients based on inference similarity, addressing the non-IID data distribution problem and enhancing the overall performance and efficiency of FL models.
Another approach for addressing the changes in clients’ input space is domain adaptation, which addresses the challenge of adapting a model trained on one domain (source domain) to perform well on a different but related domain (target domain) [113]. [114] provides a comprehensive overview of the methods and techniques used in multi-source domain adaptation (MSDA), addressing challenges such as class distribution alignment and domain shifts. There are three different approaches for domain adaptation, dissimilarity methods, sample reweighting, and Generative Adversarial Networks (GANs) [115, 116, 117]. Dissimilarity methods aim to reduce the difference between the source and target domain distributions by explicitly measuring and minimizing discrepancies. These methods often focus on aligning the statistical properties of the two domains. For instance, [118] proposes the Manifold Maximum Mean Discrepancy (M3D) to measure and minimize local distribution discrepancies. On the other hand, [119] introduces a novel method for unsupervised domain adaptation (UDA) that leverages similarity learning to improve the classification performance on an unlabeled target domain using features learned from a labeled source domain. Sample Reweighting is another approach, which involves adjusting the weights of source domain samples to make the source distribution more similar to the target distribution. In particular, [120] proposes a reweighting method for model aggregation in FL that considers the volume and variance of local datasets among clients to address the issue of class label heterogeneity. Also, [121, 122] propose methods to learn across various domains in Natural Language Processing (NLP) tasks by leveraging confidence-weighted parameter combinations, parameter combinations from multiple classifiers. Besides, GANs are used in domain adaptation to generate synthetic data that bridges the gap between the source and target domains. By training a generator and discriminator[117], GANs can create data that is indistinguishable from the target domain, facilitating better adaptation. Accordingly, several papers have contributed significant insights and methodologies, for instance, by leveraging multiple adversarial networks [123], considering a more realistic setting compared to traditional domain adaptation, where it is often assumed that the label spaces of the source and target domains are fully shared [124], using deep models, particularly when only a few labeled samples are available in the target domain [125], and addressing challenges of improving generalization in FL environments, where domain shifts occur [126]. After exploring changes in the input space and the methods for addressing these challenges, it is time to address the changes in the behavior of clients’ datasets, which are categorized into four classes of group level personalization methods, contextual information methods, federated multi-task learning, and cohort-based FL [9]. Group-level personalization effectively addresses the challenges of data heterogeneity and behavioral changes in FL, which is clarified in [127, 128]. By clustering clients based on data similarities and creating specialized models for each group, this approach ensures that models are both accurate and adaptive to changes in client behavior. For instance, [129] utilizes hierarchical clustering for local updates in FL to separate clients by the similarity of their updates, which allows for more efficient training on non-IID data and achieving higher accuracy. Moreover, contextual information methods involve techniques such as abnormal client behavior detection [130], regulated client participation [131], and multicriteria client selection (FedMCCS) [132] to handle variations in client behavior and data distribution. These methods enhance the robustness, efficiency, and accuracy of FL by effectively managing non-IID data, client participation variability, and handling resource constraints. To explore more solutions for changes in behavior, Federated Multi-Task Learning (FMTL) has been proposed as an advanced approach within FL that aims at handling client-specific tasks, while maintaining data privacy. This approach is introduced in [133] by proposing a novel systems-aware optimization method, named MOCHA. It extends the traditional FL by allowing the training of multiple tasks simultaneously across different clients. It offers a robust solution for managing diverse client behaviors by providing personalized, efficient, and fair learning frameworks [134]. These approaches improve model accuracy and generalization, while preserving client data privacy [135, 136]. Furthermore, Cohort-Based FL (CBFL) is a specialized approach within FL that aims to enhance learning performance by grouping clients with similar data distributions into cohorts. In particular, in industrial settings, CBFL can manage skewed data by organizing clients into cohorts based on similar data distributions [137]. This approach improves model performance by aligning the training process with the specific characteristics of industrial clients. [138] proposes a method that is the integration of cohort-based Kernel Principal Component Analysis (KPCA) with multi-path service routing in FL by leveraging KPCA for effective feature extraction and dimensionality reduction, forming client cohorts for homogeneous data distribution, and optimizing communication through multi-path routing, which is a powerful approach to enhance the efficiency and performance of FL systems and addresses the challenges of data heterogeneity and communication efficiency in FL.
Imbalanced Data
Besides non-IID data, imbalanced data is another subset of statistical heterogeneity, which refers to a situation where certain classes or labels in the dataset are underrepresented compared to others. This imbalance can lead to biased model performance, favoring the majority class and neglecting the minority class. Imbalanced data is categorized into two subsets, which are unequal sample sizes and skewed data distribution. When different classes in the dataset have a vastly different number of samples causes unequal sample sizes, leading to fail in correctly classifying the minority class due to the overwhelming presence of the majority class. This can result in high accuracy for the majority class but poor performance for the minority class, which is often more critical in applications such as fraud detection, intrusion detection, or disease diagnosis. To address this challenge, techniques such as under-sampling the majority class or over-sampling the minority class are commonly used, which is explored in [139, 140, 141]. On the other hand, skewed data distribution refers to the asymmetry in the data distribution for a particular variable or feature, which means that the frequency of occurrence of some classes is significantly higher than others, leading to a disproportionate representation of classes, making it difficult for models to learn the decision boundary for the minority class. This often results in poor generalization of new data, especially for the minority class. Addressing skewed distributions involves strategies such as clustering and density-based approaches to balance the data before training [142]. The upper part of Figure 1, statistical heterogeneity, is addressed through this part, in which most of the solutions are the subsets of FL. In the following part the conceptual heterogeneity and its subsets are addressed.
3.1.2 Conceptual Heterogeneity
Conceptual heterogeneity arises from differences in data distribution, computational power, and network conditions among participating nodes, leading to challenges in achieving efficient and robust training performance. This discussion focuses on the subset of conceptual heterogeneity related to data heterogeneity, where clients have different types of data. These differences manifest in label distribution or feature distribution, as depicted in the middle part of Figure 1. When considering label distribution heterogeneity, different clients may have datasets with varying sets of labels (diverse label sets). For example, one node might have data labeled for dog breeds, while another has labels for car models, complicating the training of a unified model. To address this, [133] proposes FL along with a multi-task learning approach to enable each client to learn its specific task, while simultaneously sharing and benefiting from a common representation or model, thereby improving both individual and global learning outcomes. Besides, [143] by proposing an unsupervised domain adaptation approach aims to improve the performance of a model on a target domain (which has no labeled data) by leveraging labeled data from a source domain. Finally, [144] reviews transfer learning, in which the core idea is to transfer the representations, parameters, or knowledge from the source task to the target task, thereby enhancing the model’s performance on the target task despite differences in label distributions.
This helps in dealing with diverse label sets by accommodating varying tasks and labels. Additionally, inconsistencies in how data is labeled across different clients (inconsistent labeling) can occur. For instance, one node might label an image as ”cat”, while another node labels a similar image as ”feline”, leading to confusion during model training. [3] introduces an approach in which clients can periodically align their labeling schemes through consensus mechanisms, ensuring that the labels become more consistent over time. This can be done through FL protocols that involve occasional exchanges of labeled examples or metadata. Also, [145] addresses inconsistent labeling by aligning feature distributions between source and target domains using deep adaptation layers and domain adversarial training (label normalization and standardization). This creates a shared feature space, minimizing the impact of label discrepancies. Semi-supervised learning is another approach, which is reviewed in [146], to address inconsistent labeling by leveraging a small amount of labeled data along with a large amount of unlabeled data to improve the overall labeling quality. Feature distribution heterogeneity occurs when clients have different features available in their datasets (varying feature sets) or when the scale of features varies across nodes (feature scaling differences). For example, in a healthcare system using datasets from several hospitals, one hospital might record patient data with features such as age and weight, while another hospital records features such as blood pressure and cholesterol levels, complicating model integration. Methods such as feature alignment techniques [147], data augmentation [148], and collaborative feature selection [149] are utilized for addressing this challenge. Feature scaling differences arise if one hospital records blood pressure in mmHg, while another records it in kPa. In this case, standardization and normalization [150], Federated Preprocessing [151], and Metric Learning [152] are proposed for addressing this.
When different clients participate in training a model their differences in the components and capabilities lead to system heterogeneity. This heterogeneity can be broadly categorized into two main subsets: hardware heterogeneity and software heterogeneity. Hardware heterogeneity encompasses differences in computational power, memory and storage capacities, and network connectivity among the devices. For example, some devices might have high-end processors capable of fast computations, while others may have limited processing power, leading to synchronization issues during model training. Similarly, variations in memory and storage capacities can result in some nodes struggling with large datasets or complex models, creating bottlenecks. Diverse network connectivity also plays a role, with some devices connected via high-speed broadband and others relying on slower mobile data, affecting data transmission rates and overall system efficiency. On the other hand, software heterogeneity involves the diversity of operating systems, software versions, and programming languages or frameworks used across different nodes. Devices in the system might run on various operating systems such as Android, iOS, Windows, or Linux, requiring robust software solutions that are compatible across all platforms. Different versions of the same software or libraries can cause execution inconsistencies, necessitating careful version control and updates. Additionally, the use of various programming languages and frameworks, such as Python, Java, TensorFlow, or PyTorch, requires interoperability solutions to integrate these diverse technologies seamlessly. Addressing both hardware and software heterogeneity is crucial for the efficient and effective operation of FL and distributed systems, ensuring consistent performance and reliability across all nodes.
3.2 Resource Constraints
Collaborative machine learning, where multiple users contribute their computational power and data, is promising but faces significant challenges due to resource constraints. These constraints can limit the efficiency and effectiveness of the collaborative learning process. Table 3.5 lists datasets with structural variations and drift types, relevant for studying resource limitations. Accordingly, [153] states that heterogeneity and limited computation and communication resources are the key factors that challenge machine learning at the network edge. These constraints hinder efficiency in the application of edge servers in collaborative learning processes, most especially under edge-cloud collaborative machine learning systems. This makes achieving an optimal balance between learning performance and resource consumption challenging. Additionally, in Federated Edge Learning (FEEL), the high communication costs due to the involvement of numerous remote devices pose a significant resource constraint. Efficient algorithms that speed up convergence can help alleviate these costs, which is addressed by [154]. Moreover, collaborative machine learning techniques such as FL often suffer from low resource utilisation due to sequential computation and communication, which limits their efficiency. Leveraging pipeline parallelism can significantly improve resource utilisation and reduce training time [155]. Moreover, dynamic resource allocation and task offloading in multi-access edge computing (MEC) environments can help overcome resource constraints by optimizing the use of available resources and minimizing energy consumption, while meeting delay constraints [156].
3.3 Communication Overhead
In collaborative machine learning, multiple devices work together to train a model. This process, while beneficial, introduces significant communication overhead, which can hinder efficiency and scalability. Effective strategies to reduce communication overhead are crucial for optimizing the collaborative learning process. Several factors lead to communication overhead, including communication-efficient FL, local updates before communication, compression techniques, selective communication strategies, and decentralized learning algorithms. Frequent model updates in FL involve continual exchange of updates among devices, leading to significant communication delays. Strategies such as probabilistic device selection, quantization methods, and efficient wireless resource allocation can help reduce to communication load and improve convergence speed [157, 158, 159]. These strategies collectively enhance FL’s performance and scalability by mitigating communication overhead.
3.4 Model Stability
When multiple clients participate in the training process, it is crucial to consider the global model’s ability to maintain consistent and reliable performance as it is updated over multiple rounds with data from various distributed sources. This is the essence of model stability, which involves managing the inherent variations and challenges posed by the FL environment. Key aspects of model stability include consistent performance, convergence, and scalability. Consistent performance means that the model should show stable performance metrics across different training rounds, avoiding significant fluctuations. Convergence ensures that the model parameters and performance metrics gradually stabilize, indicating effective learning and adaptation. Scalability involves the model’s ability to maintain stability and performance as the number of participating clients increases, handling the added complexity and variability without degradation. Together, these factors contribute to a robust and reliable global model that generalizes well across diverse and distributed client data.
Several techniques can be employed to enhance model stability in FL, including Federated Averaging (FedAvg), regularization, learning rate scheduling, data augmentation, client selection strategies, and adaptive aggregation. The FedAvg algorithm, introduced by [3], enhances model stability by allowing clients to perform local updates on their data and then averaging these updates to form a global model. This method significantly reduces communication overhead and maintains robust performance, even with non-IID data distributions, contributing to consistent and reliable model training across multiple rounds.
Applying regularization techniques (e.g., L2 regularization, dropout) is another approach to prevent overfitting and ensure smoother convergence. [160] demonstrates that dynamic regularization can significantly improve the robustness and effectiveness of FL models. Learning rate scheduling, which involves dynamically adjusting the learning rate, helps ensure steady progress in model updates and prevents overshooting or slow convergence.
To address model stability, enhancing the diversity of training data on clients through data augmentation is essential. [161] proposes a FL method based on data augmentation, abbreviated as FedM-UNE, which implements the classic data augmentation technique MixUp in federated scenarios without transferring raw data. This approach aims to enhance model stability and performance by leveraging the benefits of data augmentation, while maintaining data privacy and minimizing communication overhead.
Moreover, selecting a representative subset of clients for each training round is important to balance the contributions from different data distributions, which can tackle model instability. [162] proposes a client selection strategy that considers both the computational capabilities and communication bandwidth of clients to optimize the training process. By prioritizing clients with adequate resources, the method aims to reduce training time and improve the overall efficiency and performance of FL.
Using adaptive methods for aggregating model updates, such as weighting updates based on the quality or quantity of client data, is another approach for tackling the model instability problem. [163] proposes an approach that includes adaptive aggregation methods and client selection strategies to optimize the utilization of computational and communication resources. By adapting to the varying resource constraints of edge devices, the proposed framework aims to improve the efficiency, scalability, and performance of FL.
3.5 Privacy Preservation
When different clients participate in training models, privacy preservation is essential to protect sensitive data, comply with regulatory requirements, maintain user trust, mitigate risks, and enhance data utility. By implementing robust privacy-preserving techniques, organizations can ensure the secure and responsible handling of data in distributed environments, fostering a safer and more trustworthy digital ecosystem [164]. Several techniques and frameworks have been developed to address these concerns effectively.
Homomorphic Encryption (HE)
It is a powerful cryptographic mechanism that enables computations to be performed directly on encrypted data, eliminating the need for decryption during processing. This approach ensures that data privacy is maintained throughout the computational process. In the context of FL, HE can be applied to encrypt local model parameters before they are transmitted to a centralized server. The server then aggregates these encrypted parameters without decrypting them, thereby preserving the privacy of individual data points and preventing any potential exposure of sensitive information [165, 166, 167].
Differential Privacy (DP)
It is a technique that enhances data privacy by adding controlled noise to the data or model parameters, thereby preventing the extraction of sensitive information while preserving the overall utility of the data. In FL, Local Differential Privacy (LDP) can be employed to add noise to the data before it leaves the user’s device. This method ensures that the central server, which aggregates the data from various users, cannot infer any private information from the aggregated data, effectively maintaining user privacy. For instance, [168] introduces a framework where artificial noise is added to client parameters before aggregation to ensure privacy. This method, called ”noising before model aggregation FL” (NbAFL), satisfies differential privacy by adjusting noise variance. The authors also develop a theoretical convergence bound, showing a trade-off between privacy and performance, and propose a K-client random scheduling strategy to optimize this trade-off. Evaluations confirm that the theoretical results match simulations, aiding the design of privacy-preserving FL algorithms.
Secure Multi-Party Computation (SMC)
This approach allows multiple parties to collaboratively compute a function over their inputs, while ensuring that each party’s inputs remain private. This mechanism ensures that no single party gains access to all the data, preserving the confidentiality of individual inputs. In FL frameworks, SMC can be employed to securely aggregate model updates from multiple users, thereby preventing the exposure of their individual data. This method provides a robust guarantee of data privacy, making it a vital tool for maintaining security and trust in collaborative machine learning environments [169].
Table 2: List of datasets used in studies on FCL. Parameters: trial (T), channel (CH), pixel (P), and features (F). Repository: Queen Mary University of London (QMUL), Electronic Engineering and Computer Science (EECS), University of Toronto (UofT), National Institute of Standards and Technology (NIST), University of California, Irvine (UCI), Stanford University (SU), Princeton University (PU), University of New South Wales (UNSW), Canadian Institute for Cybersecurity (CIC), Ruhr University Bochum (RUB), University of Minnesota (UMN), California Institute of Technology (Caltech), University of New Brunswick (UNB), University of Extremadura (UEx), Northwestern Polytechnical University (NWPU), University of Edinburgh (UoE), Stanford University (SU), Xidian University (XDU), University of Minnesota (UMN), University of Amsterdam (UvA), Zalando Research (ZR), and Littelfuse Incorporated (LFI). Drift type: gradual (G) and abrupt (A). Data type: non-tabular (N) and multivariate (M).
{NiceTabular}
lccccccccccc Dataset Repository#Instances#Features#Classes DimensionalityData TypeDrift TypeTime-series Data FormatStructureDescriptions
Real World Dataset
CIFAR-10 [170] UofT 60K307210 32P x 32P x 3CHN✗✗ImageSImage Classification
CIFAR-100 [170] UofT 60K3072100 32P x 32P x 3CHN✗✗ImageSImage Classification
MNIST [171]NIST 70K7841028P × 28PN✗✗ImageSDigit Classification
Fashion-MNIST [172]ZR 70K7841028P × 28PN✗✗ImageSClothing Classification
EMNIST [173]NIST 2M7846228P × 28PN✗✗ImageSLetters/Digit Classification
SVHN [174]SU 630,38830721032P × 32P × 3CHN✗✗ImageSDigit Recognition
GTSRB [175]RUB 50K30724332P × 32P × 3CHN✗✗ImageSTraffic Sign Recognition
FitRec [176]UMN 102/168/253K—M✗✓TabularSRecommendation
AdultUCI 4884214214FM✗✗TabularSClassification
RCV [177]Reuters80441447236 103 -N✗✗TextSNews Classification
Caltech-256 Caltech 30,607*256* × * ×3CHN✗✗ImageS Object Recognition
ImageNet [178]SU/PU 14M19660821K256P x 256P x 3CHN✗✗ImageSVisual Object Recognition
TinyImageNet [179]SU/PU 120K1228820064P x 64P x 3CHN✗✗ImageSVisual Object Recognition
MiniImageNet✗60K2116810084P × 84P × 3CHN✗✗ImageSFew-shot Classification
ToN-IoT [3]UNSW 2.3M441044FM✗✗TabularSIntrusion Detection
CIC-IDS2017 [180]CIC 3M801580FM✗✗TabularSIntrusion Detection
AloiUvA 110250829441K192Px144Px3CHN✗✗ImageSObject Detection
HelenaUCI 6519628228FM✗✗TabularSBinary Classification
Jannis UCI 8373354454FM✗✗TabularSPerformance Prediction
NWPU-RESIS45NWPU 31500 196,608 45 256P×256P×3CH M✗ ✗ Image S Sensing Classification
CINIC-10 [181] UoE 270K 3,072 10 32P×32P×3CH N✗ ✗ Image SImage Classification
Twitter SentimentSU 1.6M636FM✗✗TextSSentiment Analysis
MLRSNet [182]XDU 109161196,60846 256P x 256P x 3CH M✗ ✗ Image SSensing Classification
Movielens [183]UMN 0.1/1/20M6/8F56/8M✗✗Tabular/TextSMovie Ratings
Sensorless Drive DiagnosisLFI/UCI 58,509481148FN✗✗TabularSMotor Fault Diagnosis
FaceScrub [184]Exposing.ai 105830128530* × * ×3CHN✗✗ImageUFace recognition
SocNav1 [185]UEx 1200 6F ✗ 6F M✗ ✗ Tabular S Social Navigation
BoTIoTUNSW 72M 46 4 46F M✗ ✗ Tabular S Intrusion Detection
NSL-KDD [186]UNB14851742542FM✗✗Tabular SIntrusion Detection
Forest Covertype [187]UCI 58101254754FM✗✗TabularS Classification
DEAP [188] QMUL-EECS 806444440T × 40CHM✗✓TabularSEmotion Analysis
Synthetic Dataset
SINE [189, 190, 191]✗ *2 2 2F N A ✓TabularS Binary classification
Circles [189, 190, 191]✗*2 2 2FNG✗TabularS Binary classification
Secure Aggregation
Secure aggregation protocols ensure that individual updates from local models are securely aggregated without exposing individual contributions. These protocols are designed so that the central server has access only to the aggregated model parameters, not the individual updates from each participant. This approach preserves user privacy by preventing the central server or any other party from accessing or inferring the private data of individual users [192]. Such protocols are crucial in FL frameworks, where maintaining the confidentiality of user data, while still enabling collaborative model training.
4 Challenges of Continual Learning
Despite the advantages of CL, streaming data inherently leads to concept drift, a phenomenon where the underlying data distribution changes over time, which can lead to a decrease in model performance as the data distribution evolves.
4.1 Concept Drift
Concept drift can be categorized into three main types: virtual drift, real drift, and hybrid drift. Table 3.5 provides examples of datasets designed to observe gradual and abrupt concept drifts.
4.1.1 Virtual Drift
Virtual drift, also known as temporary drift [193], sampling shift [194], and feature change [195], refers to changes in the distribution of input data $P(X)$ over time without altering the underlying relationship between inputs and outputs. For instance, consider an e-commerce website, where the traffic sources initially consist of 50% from search engines, 30% from social media, 10% from email campaigns, and 10% from direct visits.
Over time, due to new marketing strategies, this distribution shifts to 30% from search engines, 50% from social media, 10% from email campaigns, and 10% from direct visits. This shift, while not changing the relationship between traffic and sales, can cause the predictive model based on the original data to perform poorly. To maintain accuracy, the model must be updated to reflect the new distribution of traffic sources.
4.1.2 Real Drift
Figure 2: Taxonomy of approaches used to address global forgetting. While this taxonomy is created based on the literature on CL, it can be applied to FCL, as it also uses the principles of CL.
On the other hand, real drift occurs when the relationship between the input data and the target variable $P(y|X)$ changes over time. This means the underlying concept that the model is trying to learn is itself changing. For instance, in a financial market prediction model, the factors that influence stock prices may change due to new regulations, economic policies, or shifts in market dynamics. As a result, the relationship between the input features (e.g., economic indicators) and the target variable (e.g., stock prices) changes.
4.1.3 Hybrid Drift
Hybrid drift involves both virtual and real concept drifts occurring simultaneously. This means that there are changes in the data distribution $P(X)$ as well as changes in the underlying relationship between the features and the target variable $P(y|X)$. Hybrid concept drift is particularly challenging because it requires models to adapt to both types of changes simultaneously. For example, in a recommendation system for an online store, the types of products offered (input distribution) might change, and at the same time, customer preferences (relationship between inputs and outputs) might also change.
After exploring the various changes in data distribution caused by streaming data, this section proposes and addresses the different challenges that occur in non-stationary environments, including dealing with limited data, managing class incremental learning, and mitigating catastrophic forgetting.
4.2 Solutions
To effectively address these changes over time, Figure 2 presents a comprehensive taxonomy of concept drift and outlines solutions for each category. To address virtual concept drift, memory-based methods are proposed to help models retain the knowledge of previously seen data to adapt to changes over time effectively [196]. These methods store a subset of past data and periodically retrain the model using this data. These methods can be categorized into: rehearsal methods and generative methods. On the one hand rehearsal methods store a subset of previous data and mix it with new data during training, which can prevent catastrophic forgetting. For instance, [197] explores the use of small episodic memory buffers to mitigate catastrophic forgetting in CL scenarios. Also, a method called Maximal Interfered Retrieval (MIR) is proposed in [198], which selectively rehearses past experiences that are most likely to suffer interference from new data. Accordingly, [199] utilizes gradient information to select training samples with the most significant impact on model parameters. On the other hand, instead of storing actual data, generative methods generate synthetic data that mimics past data distributions, which helps in reducing storage requirements. For instance, [200] utilizes meta-learning to enable the model to learn continually. Inspired by human concept learning, a generative classifier is developed that effectively leverages data-driven experiences to learn new concepts from just a few samples, while remaining immune to forgetting. This approach enhances the model’s adaptability and robustness in CL environments, ensuring sustained performance and knowledge retention over time. [201] proposes an approach, which involves training a generative model to produce samples that mimic previously seen data, which are then used alongside new data to update the main learning model. This technique allows the model to maintain performance on old tasks, while learning new ones, effectively enabling CL without significant performance degradation. It is important that previously learned input domains of data are not forgotten. Regularization methods are proposed to address virtual drift by imposing constraints on the weight updates. For instance, the method in [202] focuses on dynamically adjusting the importance of features based on their relevance to the current data distribution.
To address the changes in the distribution of target data (real drift), contextual information methods and architecture-based methods are proposed. Contextual information methods utilize additional data, often external to the main dataset, to provide a deeper understanding of the context in which the data was generated. This can help in identifying patterns and changes in data distribution over time. [203] leverages contextual information by implementing a hard attention mechanism to focus on task-relevant information, thereby improving the model’s ability to retain previous knowledge, while learning new tasks without the loss of crucial previously acquired information. Additionally, architecture-based methods address changes in the underlying data distribution (real concept drift) by modifying the model’s structure to adapt to these changes effectively. For instance, [204] adapts a single fixed neural network to multiple tasks by learning binary masks that selectively activate weights, thus allowing the network to maintain high performance on both new and previously learned tasks without suffering from catastrophic forgetting. This method incurs minimal overhead and is effective for various classification tasks, even those with significant domain shifts. [205] addresses real concept drift by iterative pruning and retraining weights in a neural network, allowing it to accommodate new tasks, while preserving performance on prior tasks. This method compresses the model, freeing up capacity for new tasks and ensuring CL without performance degradation. In the next subsection, the requirements for datasets to observe different types of drifts are discussed and explored.
4.2.1 Datasets
Evaluating FCL systems is a complex challenge that requires careful consideration, since datasets need to align with the various challenges of different learning scenarios, e.g., incremental, task-specific, and domain-specific learning. For instance, Table 1 lists different dataset characteristics, including task types such as instance incremental and class incremental learning, and research domains like healthcare, cybersecurity, and IoT. This table additionally highlights specific dataset features such as data format and output types that add significant value to effective evaluation. By categorizing datasets based on these metrics, researchers are able to build a better comprehension of such challenges in performing realistic and standard evaluation of FCL systems.
To be able to use these techniques in contributive systems, our data requires some features to havewhich can be grouped into three categories. Most of the FCL research uses different publicly available datasets across various domains to validate any particular learning strategy for their challenges, including concept drift and catastrophic forgetting. Table 3.5 summarizes a comprehensive list of datasets and repositories on the basis of dimensionality, data format, type of drift, and respective repositories. The table also summarizes datasets to observe gradual and abrupt drifts with their dimensional parameters in terms of pixels, channels, or features. By providing this in-depth overview, Table 3.5 acts as a foundational point of reference for the selection of datasets that correspond to certain FCL research objectives.
The first category relates to situations, where the initial dataset of clients exhibits different distributions in behavior or both behavior and input space, which can occur in FL. Additionally, this applies if virtual drift happens or if there is no drift related to stream data. In this situation, a sufficient amount of labeled data from all participants is required at the beginning of the training process to determine their behavior. Once their behavior is established, it cannot change, so no additional labeled data is needed. On the contrary, the clients are allowed to present different conditional probabilities, but these probabilities will remain constant over time. In this situation, enough labeled data from all participants is required at the beginning of the training process to determine their behavior. Once their behavior is settled, it is not possible for it to change, so no more labeled data is required.
If the clients’ behavior changes over time but remains consistent among the devices, then sufficient labeled data from at least one participant is required periodically. Knowing the behavior of one participant is sufficient since, in these scenarios, the behaviors of all other participants will be identical. When a real concept drift occurs, the labeled data from that client will enable the model to detect the drift and respond appropriately.
5 Challenges of Federated Continual Learning
Figure 3: Fishbone diagram showing the taxonomy of challenges in FCL. Main branches show the main categories of these challenges, and sub-branches indicate specific reasons for these challenges within the specified domain.
Table 3: Main FCL approaches. Base models: Convolutional Neural Network (CNN), Recurrent Neural Network (RNN), Multi-Layer Perceptron (MLP), Reduced MobileNet-V2 (RMNv2) [206], Graph Convolutional Network (GCN), Deep Graph Library (DGL), Deep Deterministic Policy Gradient (DDPG), Fully Connected Layers (FCLs), Model-Agnostic Meta-Learning (MAML), Attention-Based Model (Att), and Vision Transformer (ViT). Aggregators: Elastic Averaging SGD (EASGD) [207], Prototypes Aggregator (PA), Weighted Ensemble-Based Aggregation (WEA), projection in turn (PIT), gradient information-based aggregation (GIA), Digital Twin Network (DTN), Gradient Integrator (GI), Batch-oriented Aggregation (BoA), Coordinate-based Model Aggregation (CooMA), and Adaptive Aggregation (AdapA). Learning strategies: supervised (S), unsupervised (U), and semi-supervised (SS). Updating: synchronous (Sync) and asynchronous (Async). Client participation: static (Stc) and dynamic (Dyn). FL architecture: centralized (Cent) and decentralized (Dcent). Streaming strategies: sliding window (SW), time window (TW), convolution kernel (CK), coreset selection(CS), adaptive windowing (ADWIN), reservoir sampling (RS), and data buffer (DB). FL Challenges: communication overhead (CO), model stability (MS), heterogeneity (H), resource constraints (RC), and privacy preservation (PP). CL challenges: concept drift (CD), class incremental learning (CI), limited incremental data (LD), and catastrophic forgetting (CF). FCL challenges: global forgetting (GF), global model overfitting (GMO), negative knowledge transfer (NT), significant accuracy loss (AL), and local forgetting (LF). * and ✗ denote not applicable and not mentioned.
{NiceTabular}
lccccccccccc Topology/Techniques Challenges
Approaches Model Homogeneity Base Models Client Participation
FL Architecture Updating Aggregator Learning Strategy Stream Handling
FLCLFCL
Fed-ReMECS [28] ✓FFNNStcCentSyncFedavgSSWPP**
ISCN [29] ✓*Stc Cent*ISCN SSW PPLD*
[30] ✓SE+ATTStc Cent*FedAvg SCK H**
FedSpace [8] ✓ResNetStc CentAsyncFedAvg S*MSCLGF
FedConD [31] **Dyn CentAsyncFedConD S* CO/MSCD*
ASO-Fed [32] **DynCent AsyncASO-FedS *H/CO CD*
[33] BGD/SGD /mini-batch SGD DynCentAsyncFedAvg SstrSAGA[208] MS/RC**
CHFL [7] **Stc CentSyncFedAvg/VFL S*MS/COCDLF
Cross-FCL [34] ✓*Dyn/Stc CentSync FedAvg/ FedAvg+EWC S Regularization Term H/COCDNT
FedDrift-Eager/FedDrift [35] ✓*Stc DcentSyncFedAvg SSW HCD*
Master-FL [36] **Stc CentSyncFedAvg/FedOMD[209] S* HCD*
SOFed [37] BYOL/ SimCLR/MoCo Dyn CentSyncFedAVGU CS HCFLF
FLARE [38] ✓CNNStc CentSyncFedAVGS TW COCD*
[39] ✓CNNStc CentSyncFedAVGS * PPCFGF/LF
FairFedDrift [40] ✓CNNStc CentSyncFedAvgS * MSCD*
[41] ✓CNNDyn CentSyncFedAvg S* CO/RC CD*
ADMM-FedMeta [42] ✓CNNStc CentAsyncInexact-ADMM S* MS/RCLD/CFLF
FedINC [43] ✓CNN/ResNetDyn CentSync FedAvg+PA S* HCD/LD/CFGF/LF
Adaptive-FedAVG [44] ✓MLP/CNNStc CentSyncAdaptive-FedAVG S* HCD/CI*
AFAFed [11] IoTNet [210] /RMNv2 Stc CentAsyncEASGD S* CO/H/RC**
O-GFML/PSO-GFML [45] ✗*Dyn DcentSyncFedAVG S*CO/RCCD*
GLFC [46] ✓ResNet [211]Dyn CentSync* S* ppCI/CFGF/LF
[47] ✓*Stc CentSyncFedAVG STW CO/RCCD*
CoOptFL/DYNAMITE [48] ✓SVM/CNNStc CentSync/Async* SRS H/RCCD*
Fed-IWDS [49] ✓CNNDyn CentAsyncFedAvgS* H/RC/PPCD*
[50] ✓CNNStc CentSyncS**
FedSKF [51] ✓ResNet/VGGStc CentSyncPITS* H/PP CI/CF GF/LF
[52] ✓ResNet/CNNStc CentAsyncFedAvg [3]S* RCCD*
[53] ✓CNNStc CentSyncFedProx [212]S* PP**
FCIL-MSN [54] ✓ResNetDyn CentSyncGIAS* CO/MS/HCD/CFGF/LF
FedADC [55] ✓CNN* CentSyncSLOWMO [213]S* MS/HCD*
FCL4SR [56] ✓ GCN+DGLDyn DcentSyncElastic Transfer S* PPCD/CFGF/NT/LF
[57] ✓CNN [214]Stc CentSyncFedAvgS* RCCFLF
FCL-SBLS [58] ✓DDPG-based algorithmDyn CentSyncDTNS* PPCFGF/LF
FedStream [59] **Stc CentSync/AsyncFedAvgSDB CO/H/RCCD*
FedStream [60] ✓K-Means+KNN [215]Stc CentSyncFedAvgS* CO/PPCD/CI*
[61] ✓CNN+FCLsStc CentSyncFedAvgS* PPCI/CFGF/LF
FedProK [62] ✓ResNetStc CentSync/AsyncFedAvgSPKF/FT HCI/CFGF
FedKNOW [63] ✓ResNet/CNNStc CentSyncFedAvgS* COCFNT/LF
FedRCIL [64] ✓ResNetStc CentSyncFedAvgSDB H/CO CF GF
LCFIL [65] ✓MLP/LeNet/VGG/AlexNetDynCentSyncS* MS
Re-Fed [66] ✓ResNetDynCentSyncFedAvgS* H CFGF/LF
ICMFed [67]✓DenseNet/EfficientNetDynCentSyncICMFed+FOMAML/ICMFed+ReptilS* CO/HCD/CFLF
LGA [68]✓ResNetDynCentSync FedAvgSDB HCI/CFGF/LF
[69]** StcCentSyncFedAvgSFIFO/SRSR/DRSRMS/RCCD*
Flash [70] ✓FCN/LSTM/CNN/ResNetDynCentSyncFlashS* HCD*
[71] ✓BERT-base model [216]DynCentSyncFedAvgSTW RC CF LF
FedPC [72]✓ResNetStcDcent✗✗S*CO/RC/PPCD/CFLF
cTD-$\alpha$MAML [73]✓MAMLStcCentSyncFedAvgS* CO/RC/PPCD/CFLF
SFLEDS [74] ✓KNNStcCentSyncKNN+Selected PrototypesSSDB PPCD/LD*
FedNN [75] ✓LeNet [171]/ResNetStcCentSyncFedDC+FedNNS*CO/MS/HCD*
Fed-SOINN [76] ✓AttDynCentSyncFedAvgSS*CO/MS**
RRA-FFSCIL [77] ✓ResNetDynCentSync*S*HLD/CFGF/LF
[78] ✓VPT360 [217]StcCentSyncFedAvgS* H/PPCFLF
[79]✓Multi-head AlexNetStcCentSyncCooMAS* H/RC CF GF/LF
[80] ✓LSTM+FCStcCentSyncAdapASSW HCD/CFGF/LF
FL-IIDS [81] ✓CNN-GRUDynCentSyncFedAvgS*MS/RCCD/CF GF/LF
FedViT [82]✓ViTStcCentSyncFedAvg/FedKNOWS*CO/MSCD/CFGF/NT/AL
CFeD [83]✓TextCNN [218]/Resnet+FCDynCentSyncCFeDS**CF GF/LF
FedET [84]✓ResNet/ViT-Base DynCentSync*S*COCI/CFGF/LF
FedWeIT [85] ✓LeNet/ResNetStcCentSync EWC methodS* COCFNT/LF
[86]✓TinyBERT [219]DynCentSyncFedAvgS*RCCFLF
[87]✓4-Conv/ResnetStcCentSyncFedAvgS*CO/RC CF LF
[88]✓ Conv+FC+LSTMStcCentSyncFedAvgS*PP**
SLMFed [89]✓CNNDynCentSyncSLMFedS*HCFLF
TARGET [90]✓ResNetDynCentSyncTargetS*PPLD/CFLF/GF
HFIN [91]✓1D-CNNStcDcent✗✗S**CFLF
[92]*VariedStcCentAsyncVariedS*HLD*
In an FCL environment, numerous devices send their updates to a global model, necessitating continuous updates due to incoming data. Devices incrementally send the learned information to the global model, introducing specific challenges unique to the FCL environment. These challenges occur locally, globally, and during the transfer of knowledge between the global and local models. Managing incoming data locally, while keeping the global model updated presents significant difficulties, underscoring the importance of addressing these global model challenges.
As illustrated in Figure 3, FCL challenges are categorized into three main areas: local challenges, global challenges, and challenges related to knowledge transfer between global and local models. Each of these categories is further divided into sub-challenges, which are discussed in the following sections. Their origins (root causes) and the proposed solutions are examined in detail, emphasizing the complexity and importance of overcoming these obstacles to achieve effective FCL.
5.1 Global Model Challenges
One of the main concerns in FCL is how clients can handle continual data, while simultaneously learning in a distributed manner. It is critical to understand how the challenges faced by clients affect global training and how these challenges can be tackled. In this section, four main challenges are discussed: significant accuracy loss, catastrophic forgetting, global model overfitting, and negative knowledge transfer. The solutions to these challenges are also explained. Table 5 summarizes approaches to address global model overfitting and catastrophic forgetting in FCL.
5.1.1 Catastrophic Forgetting
Catastrophic forgetting, also known as global forgetting, is a significant challenge in FCL, caused by class imbalance and data heterogeneity among clients in the FL environment, and class-incremental and limited incremental data issue in the CL environment. This leads the global model to adapt to dominant classes or new data distributions, often forgetting earlier tasks or less frequent classes. Also, training with a small amount of labeled data not only can deteriorate the model performance on new tasks, but also exacerbate catastrophic forgetting of previous knowledge. To address this, strategies are needed to preserve and transfer new client information despite limited resources and communication bandwidth in FL environments. Tables 1 and 3.5 list datasets supporting class-incremental and instance-incremental tasks relevant to these strategies. As depicted in Figure 2, three primary methods to mitigate global forgetting are preserving the global model, learning from exemplars, and a combination of these two.
Preserving The Global Model
To preserve the global model in FCL, knowledge fusion is used to leverage the previously trained global model, integrating new and old data effectively. Five primary methods are employed for this purpose: proxy server methods, prototypical learning, optimization and Loss function modification, knowledge distillation, and gradient-based methods.
Proxy server methods utilize intermediary servers to store and manage previous global model data. For instance, [68] selects the best old model for global anti-forgetting and augments new prototype images via self-supervised prototype augmentation. Similarly, [46] proposes a proxy server that selects the best old global model to assist local relation distillation.
Prototypical Learning focuses on representative examples to maintain class prototypes. For example, [62] uses prototype vectors to average client model representations, preserving knowledge across tasks.
Adjusting optimization techniques and loss functions balance new and old knowledge. This includes loss function modifications, learning rate adjustments, and proximal terms. [8] introduces a specific loss function to distill knowledge from the old global model, employing augmented old global prototypes and a contrastive representation loss to properly allocate new global class prototypes and align new local classes with their previous positions. Yuan et al. [72] mitigate catastrophic forgetting by adjusting the learning rate over time.
Knowledge distillation is a widely used technique for fusing knowledge in FCL [220]. The core idea involves transferring the knowledge from an already trained teacher model to a student model. Essentially, the teacher model guides the student to replicate the same mapping relationship by providing the label $Y$ for the sample $X$, similar to the way humans teach. The continual distillation mechanism uses distillation techniques at multiple scales and employs contrastive learning to transfer knowledge from old tasks to new tasks while preserving the learned knowledge [64]. Federated Not-True Distillation (FedNTD) [221] focuses on preserving the global model’s knowledge by only considering not-true classes during distillation, thus maintaining the global perspective without directly accessing local data. The Enhancer distillation method [84] incorporates a small module, the enhancer, to absorb and communicate new knowledge, thereby mitigating the imbalance between old and new knowledge. Lastly, exemplar-free distillation [81] avoids the use of a rehearsal buffer by employing distillation methods that do not require storing past data samples.
Moreover, some papers prevent the forgetting of previously learned tasks by modifying gradients. Accordingly, [63] integrates gradients from the current task with gradients from the most dissimilar previous tasks, as similar as [68], which adjusts gradient compensation dynamically to balance the forgetting rates of different categories, ensuring that no single category is disproportionately forgotten.
Learning From Exemplars
In contrast to the first approach, in which the previously learned information is preserved by fusing the previous global model and using only the data from the current task, an exemplar refers to a sample from a previous task that is stored in a memory buffer for future training that is another approach for preserving previous information. There are different methods for selecting a small portion of previous data as an exemplar, which fall into two main sections, global exemplar and local exemplar. The global exemplar approach involves the server aggregating a subset of data from clients and redistributing it to clients during training. However, this method does not adhere to the principles of FL. On the other hand, local exemplar refers to each client retaining a subset of data from previous tasks for use in future training. On the other hand, local exemplar refers to each client retaining a subset of data from previous tasks for use in future training.
Replay-based methods refer to approximating and recovering old data distributions, which is the principle way for learning from local exemplars in FCL and are divided into two sub-directions depending on the content of replay, direct and generative replay. Direct replay, known as experience replay, involves storing and replaying actual data samples from previous tasks during the training of new tasks. Replay-based methods can be optimized using several strategies such as importance-based selection, coordinated replay, split-based, and retainer mechanisms. For instance, [66] selectively stores the most critical data samples to ensure that essential knowledge is retained. This work also considers coordination among clients for better sample selection, ensuring that important samples are replayed. Also, [79] utilizes a retainer mechanism to ensure that valuable information from previous tasks is preserved. On the other hand, generative replay (pseudo-rehearsal) uses generative models to synthesize data samples that approximate the distribution of previous tasks. In this context, [39] proposes server-side generative model training, in which the server trains a generative model in a data-free manner at the end of each task. This model synthesizes images that mimic the distribution of old tasks, which can then be used to update the global model. This helps the global model retain knowledge from all previous tasks, thus mitigating global forgetting.
These approaches help in maintaining a balance between acquiring new information and retaining previously learned knowledge, and, thus, mitigating global forgetting in FCL.
5.1.2 Global Model Overfitting
Global model overfitting in FCL is a significant challenge that occurs when the global model, which is aggregated from the local models of different clients, performs exceptionally well on the training data (often non-IID and limited to specific clients) but fails to generalize to unseen or new data. Several factors contribute to this problem: resource constraints, communication overhead, heterogeneity, which are rooted in FL, class-incremental, and limited incremental data that arise due to the unique characteristics of CL. Resource constraints, such as limited computational power and memory on client devices, force the use of simplified models or reduced training iterations, which can lead to underfitting at the local level but overfitting at the global level, when these simplified models are aggregated. Communication overhead limits the frequency of model updates and necessitates the use of compression techniques that can degrade model quality, leading to overfitting as the global model may overly adapt to the noisier, and compressed updates from clients. Heterogeneity among clients, with varying data distributions and device capabilities, causes the global model to struggle with generalization, as it may overfit to the dominant clients’ data distributions, neglecting the broader population. In CL, class-incremental tasks, where new classes are introduced over time, further complicate the global model’s ability to generalize, as it may overfit to recent classes at the expense of forgetting previous ones. Finally, limited incremental data due to the sequential nature of CL forces the model to deal with sparse data, increasing the likelihood of overfitting to these small and biased datasets rather than learning a robust and generalized representation. These interconnected challenges highlight the complexity of preventing overfitting in FCL, where solutions must address the interplay between FL’s distributed nature and CL’s dynamic task environment. Overfitting may accur in two cases, when the global model focuses on the past data and overfited on them, which causes low performance and not learning the new data. On the other hand it can be that the global model is overfitted on new data, which causes catastrophic forgetting. So, the core idea of solutions is to make a balance between learning new data and also preserving previously learnt information. These solutions are categorized into five categories, traditional methods, optimization, balancing learning, generalization, and few-shot learning [77, 222]. To enhance the robustness of a machine learning model against overfitting, several key factors should be considered. These principles are equally applicable when designing the global model in FCL systems. By carefully addressing these factors, the model’s ability to generalize to new data can be significantly improved, ensuring better performance across diverse environments and reducing the risk of overfitting.
Solutions to FCL typically require an elaborate balance of learning strategies, updating mechanisms, and aggregation techniques aimed at dealing with the unique challenges of distributed and dynamic environments. These approaches often fall under supervised, unsupervised, and semi-supervised learning, with model variants ranging from CNNs and RNNs to state-of-the-art transformers such as Vision Transformers (ViTs). Aggregators such as Elastic Averaging SGD and weighted ensemble-based methods provide additional model stability and adaptability. These are all summarized in Table 5, which explains the relationships among base models, aggregation strategies, and specific challenges faced by FCL, such as catastrophic forgetting, global model overfitting, and communication overhead. This table provides a detailed reference that facilitates indicating how various techniques meet different requirements of FCL systems.
5.1.3 Negative Knowledge Transfer
Figure 4: Categorization of techniques used to mitigate negative knowledge transfer
Negative knowledge transfer in FCL refers to the phenomenon, where knowledge acquired from one task or dataset adversely affects the model’s performance on subsequent tasks or datasets. Table 5 reports models and aggregation techniques that mitigate this issue. This issue often arises when the knowledge gained from previous tasks is either irrelevant or harmful to new tasks, leading to situations where averaging user-specific models actually degrades performance for certain subjects. In the context of FCL, where multiple clients (or edge devices) continuously learn from their local data and contribute to a global model, negative knowledge transfer can occur due to several factors, including data heterogeneity, model stability, limited incremental data, and catastrophic forgetting.
One of the primary causes of negative knowledge transfer in FCL is the heterogeneity of data across clients, a common challenge in FL. When clients have data that is not identically and independently distributed (non-IID), the resulting data distribution across clients can vary significantly. This variability can cause the global model to be influenced by irrelevant or conflicting knowledge from different clients, thereby diminishing its overall performance. Additionally, limited incremental data and issues with model stability can exacerbate negative knowledge transfer. Clients typically train their models on local data, which may be biased or not representative of the global data distribution. As a result, when the global model aggregates updates from all clients, it may overfit to these local idiosyncrasies, leading to poor generalization to new tasks or data. Furthermore, as the model learns new tasks, it may experience catastrophic forgetting, where previously learned tasks are forgotten, especially if the new tasks involve learning patterns that conflict with older ones. This form of negative transfer ultimately deteriorates the model’s performance on earlier tasks as it adapts to new ones.
As shown in Figure 4, addressing negative knowledge transfer in FCL involves strategies such as careful design of aggregation methods, regularization techniques to preserve important knowledge, and ensuring that the learning process remains robust to the differences in data distributions across clients. These strategies are illustrated in Figure 4. To address this challenge, [63, 82] proposes FedKNOW and FedViT, respectively, which are client-side solutions for mitigating negative knowledge transfer and a novel concept of signature task knowledge to extract and integrate the knowledge of signature tasks that are highly influenced by the current task. FedKNOW is designed to continuously extract and integrate knowledge from signature tasks that significantly influence current tasks by assessing the relevance of tasks, when integrating knowledge. This assessment helps in filtering out knowledge from tasks that may not align well with the current task’s objectives. Each client in FedKNOW consists of a knowledge extractor, a gradient restorer, and, crucially, a gradient integrator. During the training of a new task, the gradient integrator plays a key role in preventing catastrophic forgetting and mitigating negative knowledge transfer. It achieves this by effectively combining the signature tasks identified from past local tasks with relevant tasks from other clients, facilitated through the global model. On the other hand, FedViT introduces a task-aware Vision Transformer architecture that is capable of distinguishing between different tasks and adapting its learning process accordingly, which reduces the risk of negative knowledge transfer by ensuring that the model’s focus remains on relevant features and patterns specific to each task, rather than indiscriminately applying knowledge from previous tasks that may not be applicable. Also, this paper utilizes the selective parameter update mechanism, which carefully chooses which parts of the model should be updated based on the new task. Moreover, The model in FedViT is designed with task-specific heads, which means that different tasks can have their own dedicated output layers. This approach reduces the likelihood of negative knowledge transfer by isolating the outputs for different tasks, ensuring that the learning from one task does not adversely affect the performance of another task. [34] addresses negative knowledge transfer by enhancing collaboration among edge devices to share relevant knowledge, employing selective knowledge sharing to filter out harmful information, and maintaining task awareness to prioritize valuable knowledge. Additionally, it uses gradient adjustment mechanisms to ensure new knowledge integrates without conflicting with prior learning, thereby enabling more robust and accurate learning across distributed edge environments.
5.1.4 Significant Accuracy Loss
Many factors lead to significant accuracy loss in the global model. For instance, previuos challenges in the global model in FCL resulted in poor performance. However, it may caused by other factors such as model stability, which is rooted in FL and is one of the most prominent reasons. In general, the training procesure plays an important role in the model’s capability to learn the main information, converge, and maintain good accuracy. Model complexity, hyperparameter tuning, and aggregation methods should be considered. Also, resource constraints in FL are another issue, which doesnt let the learnt information in local models aggregated completely so it cause a loss in final accuracy. Moreover, limited incremental data is another problem rooted in CL, which leads to a loss in acuuracy.
Several factors contribute to significant accuracy loss in global models, particularly in FL and CL contexts. One of the primary challenges in FL is model stability, which can lead to unacceptable accuracy in the global model. This instability often stems from the distributed nature of FL, where the training process plays a crucial role in the model’s ability to learn essential information, achieve convergence, and maintain high accuracy.
Moreover, model complexity, hyperparameter tuning, and aggregation methods are critical factors that influence the model’s performance. In FL, resource constraints can hinder the effective aggregation of learned information from local models, further contributing to accuracy loss. Additionally, in CL, the availability of limited incremental data poses a significant challenge, leading to reduced model accuracy over time.
In summary, to enhance the accuracy of global models, it is essential to address issues related to model’s stability, optimize training procedures, and carefully consider model complexity, hyperparameter tuning, aggregation methods, and resource constraints.
5.2 Knowledge Dissemination
For streaming data, asynchronous updating might be favored to allow rapid adaptation, while synchronized updating might be preferred in static environments to ensure consistency across clients.
5.2.1 Updating Frequency
Table 4: Advantages and disadvantages of synchronous and asynchronous updating in FCL.
Updates Advantages Disadvantages
Synchronous Consistency Communication strains
Simplified aggregation Scalability issues
Lower risk of staleness
Asynchronous Flexibility Staleness of updates
Scalability Complex aggregation
Fault Tolerance
In FCL, updating frequency refers to how often updates are communicated between the clients (distributed devices or nodes) and the central server. This concept is crucial in determining the efficiency, effectiveness, and overall performance of the learning system. There are two primary approaches to updating: synchronized and asynchronous. The comparison between synchronized and asynchronized updating is crucial for understanding the trade-offs in communication efficiency, model convergence, and overall system performance. Table 4 illustrates the advantages and disadvantages of each method.
In synchronized updating, all clients (or devices) train their local models on their respective data, and then, transmit the updated parameters to a central server simultaneously after a predetermined number of training iterations or epochs. The server aggregates these updates—often through methods such as averaging—to produce a global model, which is subsequently distributed back to the clients for the next round of training.
One of the key advantages of this approach is the consistency in updates, as all clients contribute to the global model simultaneously. This leads to more stable convergence and simplifies the aggregation process since all updates are received concurrently. Additionally, by using all updates at once, there is a reduced risk of outdated model parameters negatively impacting the global model.
However, synchronized updating also has its drawbacks. It necessitates that all clients be ready at the same time, which can introduce delays, particularly if some clients have slower networks or processing capabilities—a common scenario in FCL due to the dynamic nature of incoming data. Moreover, the inherent flexibility of FCL allows for dynamic client participation, enabling the inclusion of new clients and the exclusion of existing ones as training progresses. This adaptability is crucial in real-world applications, where devices may connect and disconnect intermittently, data availability may fluctuate, or new data sources may emerge. As a result, synchronized updating can become time-consuming, as the process may experience significant delays due to the need to wait for all clients to be ready simultaneously.
In asynchronous updating, clients transmit their updates to the server independently, without coordinating with other clients. The server integrates these updates into the global model immediately upon receipt, without waiting for all clients to complete their local training.
This approach offers several advantages. Clients can send updates as soon as they are ready, reducing idle times and enhancing overall system efficiency. Additionally, asynchronous updating allows the system to scale more effectively, as it does not require synchronization across a large number of clients. The system also demonstrates greater resilience to client dropouts or delays, as it does not rely on all clients being synchronized.
However, this method presents certain challenges. Updates may be based on outdated versions of the global model, potentially introducing inconsistencies and slowing convergence. The server must also manage and integrate updates that arrive at different times, which increases complexity and may necessitate more sophisticated algorithms to ensure stability.
To compare these two approaches, some works highlight that while synchronized updating ensures consistent model updates, it can increase latency due to stricter communication requirements. This is while asynchronized updating offers greater flexibility and faster updates; however, it may lead to inconsistencies in the aggregated model. Each study compares these methods across different aspects such as accuracy, efficiency, robustness, scalability, and practicality in their respective contexts [62, 85, 2, 11].
Accordingly due to the dynamic environment of FCL, the choice between synchronized and asynchronous updating in FCL depends on the specific use case, the nature of the data distribution, and the system’s constraints. Synchronized updating may be preferable in environments where consistency and stable convergence are critical, especially with static datasets, as it ensures that all clients update the global model based on the same version, maintaining accuracy and uniformity. On the other hand, asynchronous updating might be more suitable for streaming data and highly distributed systems, where flexibility, rapid adaptation, and scalability are prioritized, allowing for frequent updates and quick responses to evolving data patterns. So due to the FCL environment, in which we can see dynamic client participation and incremental data, asynchronous updating might be favored to allow rapid adaptation.
In FCL, client participation can be categorized into two types: static and dynamic. Static client participation involves a fixed set of clients that consistently contribute to the training process over time. Dynamic client participation, on the other hand, includes a variable set of clients that can join or leave the training process at different times, allowing for more flexibility and adaptability in the learning process. These two forms of participation are compared in Table 5.
While static client participation offers advantages such as consistency, stability in model aggregation, predictable resource allocation, reduced communication overhead, and enhanced data privacy due to a stable client pool, it is not well-suited for the dynamic nature of FCL. The limitations of static client participation in FCL include limited data diversity, scalability challenges, vulnerability to client dropout or failure, potential data staleness, and a lack of flexibility. Static participation may not provide the diverse and evolving datasets needed for robust model training, making it difficult to adapt to changes in data distributions or accommodate new clients. This rigid structure hinders the continuous adaptation and real-time data updates essential for effective learning in dynamic environments. Thus, static client participation fails to meet the adaptable and scalable requirements of FCL, which thrives on dynamic client engagement and continuous data contributions to maintain model relevance and accuracy over time.
Dynamic client participation in FCL addresses the challenges of modern environments, such as client addition and removal, limited memory, communication overhead, and increased latency due to a large number of clients [33]. This approach allows the system to adaptively include or exclude clients during the training process, accommodating changes in client availability, connectivity, and data quality.
Dynamic participation offers several advantages, including enhanced adaptability to changing conditions, efficient resource utilization, improved privacy, and better model generalization. Although this approach introduces challenges such as increased complexity in managing clients, communication overhead, potential data imbalance, security risks, and latency issues, these can be mitigated through the use of appropriate algorithms and robust coordination mechanisms. The ability to dynamically adjust to real-world conditions and continuously enhance the model makes dynamic client participation an effective and flexible strategy for FCL in diverse and real-world applications.
Table 5: Summary of advantages and disadvantages of static and dynamic client participation.
Advantages Disadvantages
Static Consistency and stability
Simplified model aggregation
Predictable resource allocation
Lower communication overhead
Data privacy and security
Limited data diversity
Scalability challenges
Client dropout or failure
Data staleness
Lack of flexibility
Dynamic Improved adaptability
Efficient resource utilization
Enhanced privacy and security
Mitigation of stragglers
Robust in real-world scenarios
Improved generalization
Continuous model improvement
Complex client management
Communication overhead
Data imbalance and bias
Security risks
Computational overhead
Latency issues
Dependent on accurate data
Complex coordination
5.3 Client Models Challenges
Figure 5: Taxonomy of methods used to tackle local forgetting.
Local catastrophic forgetting refers to the phenomenon where an artificial neural network loses previously acquired knowledge about specific parts of the input space when it learns new information. This issue is particularly prevalent in sequential learning tasks, where the model updates its weights based on new data, causing interference with the knowledge acquired from earlier data. Unlike global catastrophic forgetting, which affects the overall performance on earlier tasks, local catastrophic forgetting impacts specific local regions or features within the learned tasks.
The significance of local catastrophic forgetting is profound in real-world applications, where continuous learning is required, such as autonomous driving, medical diagnosis, and robotics. Understanding and mitigating this problem is crucial for the development of robust and reliable AI systems capable of lifelong learning. As shown in Figure 5, researchers have explored various strategies to address local catastrophic forgetting, including regularization techniques, memory replay methods, and architectural modifications to neural networks.
Client catastrophic forgetting is another FCL challenge that has roots in both FL and CL challenges. If clients had sufficient memory and no constraints, this challenge could be mitigated. However, devices often lack adequate memory, which is a fundamental issue in FL. Additionally, when incoming data exhibits significant distribution changes (drift) or introduces new tasks (class-incremental learning), it becomes difficult for client models to retain previously learned information, while acquiring new knowledge. When incoming data is not enough, the model has more difficulty in forming strong representations of new knowledge and is, therefore, shallow and forgetful. Under this condition, memory constraints lead to overfitting, as the model ends up learning the limited available dataset instead of proper generalization. In addition, this disparity decreases the model’s memory capacity for new knowledge, especially in non-IID cases, where the data is less representative of distribution shifts or new tasks.
By identifying and addressing local catastrophic forgetting, we can enhance the adaptability and efficiency of AI systems, ensuring that they retain valuable knowledge over time, while continuously learning from new experiences. The solutions are categorized into distillation-based methods, replay-based methods, rehearsal-based methods, optimization algorithms, and gradient-based methods.
Distillation-based methods involve transferring knowledge from an existing model (the teacher) to a new model (the student) in a way that helps in preserving previously learned information while accommodating new data [57, 81]. [64] extracted meaningful representations and performed continual distillation locally. [84, 73] utilized enhancer and replay mechanisms to balance old and new knowledge and enhance learning.
Many studies explore replay-based methods, which can be divided into direct replay and generative replay [201, 54]. These methods store a subset of past experiences or data and periodically replay them to the model, while learning new information, helping to maintain the model’s performance on previously learned tasks. Direct replay involves storing actual data samples from previous tasks and interleaving these samples with new data during the training of new tasks [54, 71, 223]. Some studies [37, 86, 66] select the most representative samples for the replay buffer. On the other hand, generative replay involves using generative models to synthesize data samples that approximate the distribution of previous tasks. Instead of storing actual data, the method generates representative samples from the learned distribution of past tasks during the training of new tasks [39, 224]. Rehearsal-based methods differ in that they generate or synthesize data representations (such as embeddings or features) to approximate the data distribution of previous tasks. Prototypical learning [87], progressive neural networks [7], and attention models [76] are used in FCL frameworks to preserve these embeddings according to their importance [64] and use them for retaining previously learned information.
Gradient-based methods aim to mitigate client forgetting by manipulating the gradient during local training in each client separately to ensure that updates made, while learning new tasks, do not interfere with the knowledge of previously learned tasks. For instance, Dong et al. [68, 46] introduce gradient compensation loss to ensure that learning new classes does not lead to forgetting old classes. Additionally, orthogonal gradient aggregation [225] ensures that updates for new tasks are orthogonal to those for old tasks by minimizing the overlap between gradient directions for different tasks.
Some papers propose optimization and loss function modifications to tackle local forgetting [5]. Split-based methods, knowledge extraction, and local optimization are used as optimization methods. For example, specific modules designed on the client side propose loss functions such as class-projection relation strengthening loss function, class-aware distribution balancing loss function [77], and proximal term loss [72] to alleviate forgetting from streaming data.
6 Furture Research Directions
A pressing concern in FCL is the effective management of heterogeneity, which manifests both in the statistical properties of client data and in system resources. Variability in data distributions across clients poses significant challenges, particularly in non-IID scenarios. To address this, future research should focus on dynamic domain adaptation techniques that align data distributions across clients over time, enabling seamless collaboration. Personalization remains another key area of exploration. While global models aim for generalization, client-specific or group-level personalized models could balance local and global performance, enhancing adaptability and preserving the system’s overall efficiency and privacy.
Catastrophic forgetting, a major challenge in CL, becomes even more complex in federated systems, where global and local forgetting interact. Traditional methods such as regularization and replay need to be adapted for distributed and resource-constrained environments. For instance, new regularization strategies could preserve knowledge across tasks without excessive computational overhead, while exemplar-free generative replay techniques may offer memory-efficient solutions by synthesizing past data representations instead of storing them. These methods are crucial for maintaining performance on previous tasks while adapting to new ones in decentralized learning frameworks.
FCL systems inherently operate in environments with resource limitations and significant communication overhead. Efficient use of bandwidth and computational resources is critical for their scalability. Innovative parameter compression methods and sparsity-inducing algorithms can reduce communication costs without compromising model accuracy. Furthermore, asynchronous updating mechanisms, which allow clients to transmit updates independently, can enhance scalability by accommodating dynamic participation and minimizing delays. These strategies would ensure that FCL systems remain robust and responsive even in environments with fluctuating resource availability.
Privacy preservation is a cornerstone of FL and, by extension, FCL. As FCL relies on decentralized data, robust privacy mechanisms are essential. Research should continue to refine secure aggregation techniques, enabling efficient and private communication between clients and central servers. Homomorphic encryption and differential privacy approaches must be optimized for continuous learning scenarios, balancing the trade-off between privacy and model utility. The development of lightweight privacy-preserving algorithms is particularly crucial for deployment on resource-constrained edge devices.
The integration of multi-task and transfer learning paradigms into FCL represents a promising avenue for extending its capabilities. Cross-domain knowledge transfer can help models generalize better by leveraging shared patterns across tasks, especially in settings with sparse or incomplete data. This requires new methods to minimize negative knowledge transfer, ensuring that information from one task does not harm performance on another. Task-aware architectures, designed to differentiate representations for different tasks, could enhance scalability and reduce task interference, making FCL systems more effective in diverse and dynamic environments.
Handling concept drift, where data distributions change over time, is another critical area for FCL. Context-aware adaptation techniques that leverage external contextual information could allow models to detect and adjust to such changes dynamically. Similarly, incorporating incremental and few-shot learning methods can enable systems to learn effectively from minimal new data, maintaining their adaptability without sacrificing prior knowledge. These capabilities are particularly important in non-stationary environments, where data streams continuously and unpredictably.
The evaluation of FCL systems also requires significant advancements. Current benchmarks and metrics fail to fully capture the unique challenges of FCL, such as simultaneous global and local learning, dynamic client participation, and resource constraints. Holistic benchmarks that simulate real-world FCL scenarios, including non-IID data and class-incremental tasks, are essential. Moreover, new metrics should assess not only accuracy and forgetting but also fairness among clients and system efficiency, providing a comprehensive view of FCL performance.
Finally, the applicability of FCL extends across various domains, each with unique requirements. In personalized healthcare, FCL can enable continuous, privacy-preserving learning from patient data, while in IoT applications, it can support real-time analytics for highly dynamic environments. Edge-based autonomous systems, such as self-driving cars or smart drones, would particularly benefit from FCL’s ability to adapt to evolving data and tasks. Tailoring FCL methods to the specific challenges and constraints of these domains could significantly expand its impact.
7 Conclusion
This paper provides an overall survey on FCL, exploring the intersection of FL and CL for tackling dynamic, distributed, and privacy-sensitive environments. We reviewed some key challenges in FCL, such as catastrophic forgetting, concept drift, and heterogeneity of data, together with existing methodologies and solutions. The survey also brings out the role of aggregation strategies, client participation models, and learning techniques for ensuring stability and scalability in FCL systems. Furthermore, we have identified important research gaps regarding the development of more efficient resource management and privacy-preserving approaches for decentralized settings. Through this survey, we aim to provide a foundation for future works in FCL, guiding the development of adaptive, scalable, and robust models for real-world applications.
Acknowledgements
This work is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) under funding reference numbers CGSD3-569341-2022 and RGPIN-2021-02968.